OpenStudy (anonymous):

Part 1: Explain, in complete sentences, the relationship between radicals and complex numbers.

OpenStudy (campbell_st):

in complex numbers there is a convienient relationship -1 = i^2 so when working with radicals and having a negative value then the negative can be replaces by i^2 e.g. \[\sqrt{-20} = \sqrt{20i^2}\] now you can simplify the radical by removing the squares... \[\sqrt{20i^2} = \sqrt{4i^2 \times 5} = 2i \sqrt{5}\]

OpenStudy (anonymous):

Thank you very much.